How do you solve this quadratic equation?
x^2 − 16 = 0

Answer by richwmiller(17219)

Here are two ways:
1)
x^2=16
x=+-4
x=-4
x=4
x^2=16
2)
factor
(x+4)*(x-4)=0
x+4=0
x=-4
x-4=0
x=4